OpenStudy (anonymous):

Are there any holes or asymptotes in the graph of f(x)= x^2-x-6 -------- x^2-2x-3

myininaya (myininaya):

horizontal asympotes are the easier and require no math they horizontal asymptote is y=1/1=1

myininaya (myininaya):

easiest*

myininaya (myininaya):

f(x)=(x-3)(x+2) --------- (x-3)(x+1) since the x-3 cancels this implies there is a hole at x-3=0 or x=3

OpenStudy (anonymous):

\[{x^2-x-6 \over x^2-2x-3}={(x-3)(x+2) \over (x-3)(x+1)}={x+2 \over x+1}\]. So yeah there are asymptotes, one of them is the vertical asymptote at \(x=-1\).

myininaya (myininaya):

and there is a veritcal asymptote at x+1=0 or x=-1

myininaya (myininaya):

anwar!

OpenStudy (anonymous):

Hey! :D

myininaya (myininaya):

:)

OpenStudy (anonymous):

How have you been?

myininaya (myininaya):

i think i need some pills that calm me down lol how about you?

OpenStudy (anonymous):

Take some :P.. I am alright!!

myininaya (myininaya):

lol